OpenStudy (anonymous):

Find the general solution of y'+ay=0.

ganeshie8 (ganeshie8):

it is separable : \[\large \begin{array} \\ \dfrac{dy}{dx} &= - ay \\ \dfrac{dy}{y} &= -a dx\\ \int \dfrac{dy}{y} &= \int -a dx \end{array}\]

ganeshie8 (ganeshie8):

integrate ^^

OpenStudy (anonymous):

Let me try. Please don't go.

ganeshie8 (ganeshie8):

okie

OpenStudy (anonymous):

So it's ln abs(y)=-ax+C where abs means absolute value of, y=e^(-ax), right?

ganeshie8 (ganeshie8):

yep ! below form looks better : \[\large y = Ce^{-ax}\]

OpenStudy (anonymous):

Thank you!

ganeshie8 (ganeshie8):

np :)